# pythonradex: a fast Python re-implementation of RADEX with extended functionality

### v1

@@
A common task in astronomical research is to estimate the physical parameters (temperature, mass, density etc.) of a gas by using observed line emission. This often requires a calculation of how the radiation propagates via emission and absorption (“radiative transfer”). In radio and infrared astronomy, the Fortran code RADEX ([]) is a popular tool to solve the non-LTE radiative transfer of a uniform medium in a simplified geometry. I present pythonradex, a Python re-implementation of RADEX. Written in Python, it provides an easy and intuitive user interface, improved performance as well as additional functionality (continuum effects and overlapping lines) not included in RADEX.
@@
pythonradex supports four geometries: two static geometries (slab and sphere), and two large-velocity-gradient (LVG) geometries (again slab and sphere). In the LVG approximation, it is assumed that all regions of the source are Doppler shifted with respect to each other due to a velocity gradient. This means that all photons escape the source, unless absorbed locally (i.e. close to the emission location, e.g. []).
@@
To calculate line fluxes, pythonradex uses different formulae depending on the geometry (see the pythonradex documentation for more details). On the other hand, RADEX always uses the formula for a slab. Figure 3 illustrates the consequences. For a static sphere, by using the slab formula, the flux is overestimated by a factor 1.5222The factor 1.5 corresponds to the volume ratio of a “spherical slab” (i.e. a cylinder) to a sphere. in the optically thin limit. In the optically thick case, only the surface of the static sphere is visible, so the different formulae give the same result. On the other hand, for the LVG sphere, the difference is always a factor 1.5 regardless of optical depth. This is a consequence of the LVG assumption that photons always escape unless absorbed locally. By computing the optically thin flux directly [^3] In units of W/m2, $F_{\mathrm{thin}}=V_{\mathrm{sphere}}n_{2}A_{21}\Delta E\frac{1}{4\pi d^{2}}$ with $V_{\mathrm{sphere}}=\frac{4}{3}R^{3}\pi$ the volume of the sphere, $n$ the number density, $x_{2}$ the fractional level population of the upper level, $A_{21}$ the Einstein coefficient, $\Delta E$ the energy of the transition and $d$ the distance., it can be confirmed that the formulae used by pythonradex are correct.

### v2

@@
A common task in astronomical research is to estimate the physical parameters (temperature, mass, density etc.) of a gas by using observed line emission. This often requires a calculation of how the radiation propagates via emission and absorption (“radiative transfer”). In radio and infrared astronomy, the Fortran code RADEX ([]) is a popular tool to solve the non-LTE radiative transfer of a uniform medium in a simplified geometry. This paper presents pythonradex, a Python re-implementation of RADEX. Written in Python, it provides an easy and intuitive user interface, improved performance as well as additional functionality not included in RADEX, such as continuum effects and overlapping lines. In addition, pythonradex provides a self-consistent computation of the total flux for all geometries, including spherical geometries.
@@
pythonradex supports four geometries: two static geometries (slab and sphere; note that RADEX does not support a static slab), and two large-velocity-gradient (LVG) geometries (again slab and sphere). In the LVG approximation, it is assumed that all regions of the source are Doppler shifted with respect to each other due to a velocity gradient. This means that all photons escape the source, unless absorbed locally (i.e. close to the emission location, e.g. []).
@@
To calculate line fluxes (energy per time per area, e.g. W/m2), pythonradex uses different formulae depending on the geometry (see the pythonradex documentation for more details). On the other hand, RADEX always uses the formula for a slab. Figure 3 illustrates the consequences. For a static sphere, by using the slab formula, the flux is overestimated by a factor 1.5222The factor 1.5 corresponds to the volume ratio of a “spherical slab” (i.e. a cylinder) to a sphere. in the optically thin limit. In the optically thick case, only the surface of the static sphere is visible, so the different formulae give the same result. On the other hand, for the LVG sphere, the difference is always a factor 1.5 regardless of optical depth. This is a consequence of the LVG assumption that photons always escape unless absorbed locally. By computing the optically thin flux directly [^3] In units of W/m2, $F_{\mathrm{thin}}=V_{\mathrm{sphere}}n_{2}A_{21}\Delta E\frac{1}{4\pi d^{2}}$ with $V_{\mathrm{sphere}}=\frac{4}{3}R^{3}\pi$ the volume of the sphere, $n$ the number density, $x_{2}$ the fractional level population of the upper level, $A_{21}$ the Einstein coefficient, $\Delta E$ the energy of the transition and $d$ the distance., it can be confirmed that the formulae used by pythonradex are correct.